# A reanalysis of the FDA’s benefit-risk assessment of Moderna’s mRNA-1273 COVID vaccine stratified not only based on age and sex but also on prior-infection and comorbidity status

### v3

@@
Abstract. We demonstrate that the outcome of a benefit-risk assessment for vaccination may be dramatically impacted by accounting for the benefits derived from prior infection by the disease against which the vaccine is targeted. The United States Food and Drug Administration (FDA) conducted a benefit-risk assessment for Moderna’s COVID vaccine mRNA-1273 prior to its full approval, announced 1/31/2022. The FDA’s assessment focused on males of ages 18–64 years because the agency’s risk analysis was limited to vaccine-attributable myocarditis/pericarditis (VAM/P) given the excess risk among males. The FDA’s analysis concluded that vaccine benefits clearly outweighed risks, even for 18–25-year-old males (those at highest VAM/P risk). We reanalyze the FDA’s benefit-risk assessment using information available through the third week of January 2022 and focusing on 18–25-year-old males. We use the FDA’s framework but extend its model by accounting for protection derived from prior COVID infection, finer age-stratification in COVID-hospitalization rates, and incidental hospitalizations (those of patients who test positive for COVID but are being treated for something else). We also use more realistic projections of Omicron-infection rates and more accurate rates of VAM/P. With hospitalizations as the principal endpoint of the analysis (those prevented by vaccination vs. those caused by VAM/P), our model finds vaccine risks outweighed benefits for 18–25-year-old males, except in scenarios projecting implausibly high Omicron-infection prevalence. Our assessment suggests that mRNA-1273 vaccination of 18–25-year-old males generated between 16% and 63% more hospitalizations from vaccine-attributable myocarditis/pericarditis alone compared to COVID hospitalizations prevented (over a five-month period of vaccine protection assumed by the FDA). The preceding assessment derives from model inputs based on data available at the time of the FDA’s mRNA-1273 assessment. Moreover, these inputs as well as model outputs are validated by subsequently available data.
@@ 1. Introduction
Because vaccines can be given to large populations of healthy persons and can be introduced as mandatory, robust benefit-risk analyses must be conducted to ensure that vaccines have highly favorable benefit-risk profiles [1].
@@
The Food and Drug Administration (FDA) issued emergency-use authorizations for mRNA COVID-19 vaccines from Pfizer (BNT162b2) and Moderna (mRNA-1273) in December 2020 [2]. Once the mRNA COVID-19 vaccines were given to large populations, it became evident that they could cause myocarditis or pericarditis (hereafter myo/pericarditis), inflammation of the heart or of the membrane surrounding the heart, respectively, with studies indicating the highest risk among males younger than 40 years [3, 4, 5, 6, 7, 8, 9].
@@
The FDA has published two benefit-risk analyses of primary series mRNA vaccination, the first for BNT162b2 [10] and the second for mRNA-1273 [11]. The FDA presented its analyses through scenarios involving different modeling assumptions to account for uncertainties in pandemic dynamics as well as in the benefits and risks of vaccination. To assess vaccine benefits, the FDA estimated vaccine-preventable COVID-19 cases, hospitalizations, intensive-care unit (ICU) admissions, and deaths [10, 11]. To assess vaccine risks, the FDA estimated vaccine-attributable myo/pericarditis (VAM/P) cases, hospitalizations, ICU admissions, and deaths [10, 11]. The FDA concluded that the benefits of the mRNA vaccines outweighed VAM/P risk even for males in its highest-risk age-groups: 16–17 years for BNT162b2 [10] and 18–25 years for mRNA-1273 [11].
However, these FDA benefit-risk assessments omitted important inputs in the analysis of both the hospitalizations-prevented benefit of vaccination and VAM/P risk. We focus mainly on the FDA’s analysis for Moderna’s mRNA-1273 to simplify our exposition—there are significant differences in modeling assumptions in the FDA’s Pfizer and Moderna assessments. Also, VAM/P risk is substantially greater for mRNA-1273 than for BNT162b2 [3, 4, 5, 6, 7, 8, 9].
For the FDA’s assessment of mRNA-1273, benefits of vaccination were overestimated. First, the FDA did not account for the protection conferred by prior infection; and, at the time of the FDA’s Moderna analysis (mid-January 2022), likely over 70% of 18–25 year-olds in the U.S. had been infected by COVID-19—see Supplement 1. Second, the FDA did not account for incidental COVID-19 hospitalizations. Third, the FDA overestimated benefits for young males by assuming the COVID-hospitalization rate for 18–25 year-olds is equivalent to that for 26–35 year-olds and 36–45 year-olds. Finally, the FDA assumed in its most likely scenario that case and hospitalization rates would remain constant at peak 2021 Omicron levels for 5 months.
By contrast, the FDA underestimated mRNA-1273 VAM/P risks; see, e.g.,Table 3 of Section 2.3.
@@
Here, our primary aim is to demonstrate how to incorporate the protection derived from prior infection into a population-level quantitative benefit-risk model for a two-dose vaccination series. Our model has hospitalizations as its endpoint, comparing the hospitalizations-prevented benefit of vaccination to the hospitalization risk attributable to vaccination. We illustrate the utility of our model by applying it to extend and improve the FDA’s mRNA-1273 assessment relative to its hospitalizations endpoint, which can be viewed as the most important of the agency’s endpoints owing to the public-health concern that during waves of the COVID pandemic hospitals might be so overwhelmed that they would not be able to deliver high quality care to their patients.
@@
In addition to prior-infection protection, our modeling accounts for incidental hospitalizations, age-specific infection-hospitalization rates, and a more accurate assessment of VAM/P risk. We apply our model using data available to the FDA at the time of its benefit-risk analysis for mRNA-1273, assumed to be the 3rd week of January 2022 for reasons described in Section 2.1. We present evidence, using real-world data, that our model’s estimates of the hospitalizations-prevented benefit of mRNA-1273 vaccination are significantly more accurate than those the FDA obtained in [11]. We also analyze benefits and risks of vaccination for specific subpopulations such as males 18–25 having prior-infection protection and no comorbidities. Our model can be adapted to produce for any vaccine designed for deployment on a population-wide scale a benefit-risk assessment stratified not only based on age and sex but also on prior-infection and comorbidity status.
2. The FDA’s mRNA-1273 Benefit-Risk Assessment
@@
In [11], the FDA describes its benefit-risk modeling of Moderna’s mRNA-1273 for males ages 18 to 64 years. Outcomes are provided via six scenarios reflecting uncertainties in three major model inputs: COVID-19 incidence, vaccine effectiveness against cases (VE) as well as hospitalization (VEH), and vaccine-attributable myo/pericarditis (VAM/P) rate. At the time of the FDA’s mRNA-1273 analysis, the CDC was recommending a 5-month interval between primary-series completion and a booster dose. Consequently, the FDA evaluated benefits and risks over a 5-month period, which we take to be 1/1/22 – 5/31/22. The FDA assumed that all vaccine effectiveness estimates remained constant over this 5-month evaluation period. The FDA’s major modeling assumptions are summarized in Table 1 below.
Table 1. The FDA’s mRNA-1273 model-input assumptions for 18–25-year-old males [11, Table 1].
@@
Scenario COVID-19 Incidence Vaccine Effectiveness VAM/P Rate Per Million 2nd doses [11, Table 2]
@@ 6 Same as Scenario 1 Same as Scenario 1 241 cases, 207 hospitalizations
Using inputs from first row of Table 1 and incidence data from [11, Table 3], the FDA concluded for Scenario 1, its “most likely scenario,” that vaccinating one million males ages 18–25 years with two doses would prevent, over 5 months, 82,484 COVID-19 cases, 4,766 COVID-19 hospitalizations, 1,144 ICU admissions, and 51 deaths due to COVID-19, while risking 128 VAM/P cases, 110 VAM/P hospitalizations, 0 ICU admissions and 0 deaths from VAM/P [11, Table 2]. .
@@ 2.1. The Time of the FDA’s mRNA-1273 Analysis
Although the FDA does not define in [11] the time of its benefit-risk analysis for mRNA-1273, a natural choice is the 3rd week of January 2022, which concluded 10 days before the FDA’s announcement of the full approval of mRNA-1273 on 1/31/2022 [12]. This choice is consistent with the FDA’s time-of-analysis definition for its BNT162b2 assessment: the 2nd week of August 2021 [10, Section 2.1], which concluded nine days before the FDA’s announcement of the full approval of BNT162b2 on 8/23/2021 [13]. Consequently, our reanalysis of the FDA’s benefit-risk assessment [11] is based on model inputs derived from information available before 1/22/2022. Subsequent information validating our input choices will often be provided but only via footnotes. Starting in Section 4.2, we discontinue basing our modeling on information only available before 1/22/2022; in Section 4.2, our intent is to show how our model can be modified to incorporate a comorbidity (obesity) as a risk modifier.
2.2. The FDA’s mRNA-1273 Benefits Assessment
@@
where $VE$ and $VEH$ are, respectively, the effectiveness of full mRNA-1273 vaccination against cases and against hospitalization; the FDA assumed that for Omicron $VE=0.30$ and $VEH=0.72$, as noted in Table  1 above.
For Scenario 1, its most likely scenario, and two others, the FDA assumed that peak 2021 Omicron-case rates (late December 2021) would be sustained for 5 months, resulting in an exceptionally high projection of COVID-19 cases among the unvaccinated; e.g., $C=274,947$ for 18–25-year-old males. Similarly, for these scenarios, the FDA assumed that the CDC’s reported late-December 2021 Omicron hospitalization rate for unvaccinated 18–45-year-old males would be sustained for 5 months at approximately 43.54 cases per million daily [11, Table 3] and assumed that the same rate would apply to the younger 18–25 bracket, arriving at $H=6,619$ hospitalizations. Thus, for these scenarios, the FDA predicted that full mRNA-1273 vaccination of 1 million 18–25-year-old males would prevent $274,947\cdot 0.30\approx 82,484$ COVID cases and $6619\cdot 0.72\approx 4,766$ COVID hospitalizations over the 5-month evaluation period [11, Table 2].
The following factors contributed to the FDA’s overestimation of vaccine benefits:
$\bullet$
@@
The FDA did not account for the protection conferred by prior infection. The FDA’s disregarding prior-infection protection in its mRNA-1273 assessment is difficult to justify. Dan et al. found “About 95% of subjects retained immune memory at $\sim$6 months after [COVID] infection” [14]. A meta-analysis based on studies published December 2020 through August 2021 demonstrated that protection from prior COVID infection “is, at least, equivalent to the protection afforded by complete vaccination of COVID-naïve populations” [15]. The CDC acknowledged in a science brief released 10/29/21 that “The immunity provided by vaccine and prior infection are both high” [16], while another CDC study found that during the Delta wave unvaccinated persons with prior COVID infection had significantly greater protection against infection and hospitalization than did fully vaccinated, COVID-naïve persons [17]. Moreover, in December 2021, the FDA acknowledged, “The omicron variant has significantly more mutations than previous SARS-CoV-2 variants, particularly in its S-gene, the gene that encodes the virus’s spike protein”[18], suggesting that prior infection, entailing exposure to all SARS-CoV-2 proteins (such as the nucleoprotein), would likely yield greater protection against Omicron than would mRNA-1273 vaccination, which entails exposure only to the spike protein of the ancestral strain of SARS-CoV-2. Among the findings of a study of SARS-CoV-2 specific B and T cell memory is “CD8+ T cell responses preferentially target the nucleoprotein, highlighting the potential importance of including the nucleoprotein in future vaccines” [19]. [^1] Consistent with the evidence presented in this paragraph, a Qatari study found prior infection provided 52.2% effective protection against symptomatic Omicron infection and 91% effective protection against “severe, critical, or fatal” Omicron, with corresponding effectiveness estimates for 2 doses of mRNA-1273 being 2.2% and 66.3%, respectively [20, Figure 2C, 2D] (study period 12/23/21 – 2/21/22). Findings from the meta-analyses and systematic reviews [21, 22, 23, 24] provide convincing evidence that prior infection provided greater protection against Omicron during the evaluation period than did two-dose vaccination among the COVID-naïve [Supplement 1, Section 7].
@@
Further evidence of the need to account for prior-infection protection includes (i) a CDC estimate that 54.9% of 18–49 year-olds had been infected by COVID-19 by 9/30/2021 [25], (ii) a South African study [26] released 12/14/21 suggesting overall effectiveness of prior-infection protection against Omicron to be about 47% [Supplement 1, Section 4], and (iii) data from England [27] released 12/22/21 suggested that prior infection among the unvaccinated provides significant protection from Omicron hospitalization [Supplement 1, Section 5].
@@
$\bullet$ The FDA did not account for incidental COVID-19 hospitalizations. At the time of its mRNA-1273 assessment, incidental-hospitalization rates were well known to be substantial. On 1/9/22, the director of the CDC stated, “[U]p to 40 percent of the patients who are coming in with COVID-19 are coming in …with something else” [28]. [^2] We searched unsuccessfully for published CDC incidental-COVID-hospitalizations data. Figure 2 of a large U.S. study suggests an incidental rate for 0–18 year-olds typically well above 50% during its study period 3/1/2020–1/31/22 and for those 19–64 an incidental rate above 40% for 1/1/2022–1/31/2022 [29]. On 1/10/22, Ontario began publishing incidental rates, with rates for January 8–15 exceeding 45% [30]. There is evidence that incidental rates are especially high for the young. For example, a major study [31] found, “As of May 2021, the cumulative rate for hospitalization associated with SARS-CoV-2 infection was 35.9 per 10,000 children” and “When limiting the analysis to include only patients with COVID-19 who required therapeutic interventions, the hospitalization rate decreased 5.5-fold to 6.5 per 10,000 children,” which suggests an incidental rate of over 80% “for children” (persons 0-17-years-old). See Supplement 1 for additional incidental-hospitalization data.
$\bullet$ The FDA overestimated the proportion of projected COVID cases requiring hospitalization. All of the FDA’s projected case-hospitalization rates, which are described in Table 2 below, are over four times the Omicron case-hospitalization rate ($0.58\%\approx 56/9,585$) for the unvaccinated of all ages based on Omicron cases in England tested between 1 December 2021 and 14 December 202 [27, Table 3]; moreover, all are well over twice the case-hospitalization rate (0.92%) for the unvaccinated of ages 18–29 in Ontario 12/14/20 – 12/12/21 [32, Table A1a]. [^3] The FDA case-hospitalization rates in Table 2 are all well over 5 times the Omicron rate 0.42% for the unvaccinated of all ages reported in Table S10 of [33] (data from 12/15/21 – 1/17/22). Based on information available at the time of the FDA’s mRNA-1273 assessment, we expect the COVID-case hospitalization rate for the period 12/14/20 – 12/12/21 to be higher than that for Omicron; e.g., “The risk of hospital admission from emergency departments with Omicron was approximately one-third of that for Delta (Hazard Ratio 0.33, 95% CI: 0.30 to 0.37)” [34, p. 4]; see Supplement 1 for more examples. Furthermore, we expect the hospitalization rate for those 18–25 to be lower than that for 18–29 [35]. Factors contributing to the FDA’s overestimating case-hospitalization rates for the 18–25 age range include (i) assuming hospital-admission rates were uniform for ages 18–45, contradicting CDC models [36], and (ii) including incidental COVID hospitalizations. Our source for the 0.58% rate mitigated the incidental-hospitalization issue [27, p. 10], while our source [32] for the 0.92% rate did not.
Table 2. The FDA’s case-hospitalization-rate projections for unvaccinated 18–25-year-old males over the evaluation period 1/1/22 – 5/31/22 for the Omicron-based scenarios of [11, Table 2].*
Scenario Projected Cases Projected Hospitalizations Projected Case-Hospitalization Rate
11   274,947   6,619 $\frac{6619}{274,947}\approx 2.41\%$
2 89,017 2,900 $\frac{2900}{89,017}\approx 3.26\%$
3 13,010 882 $\frac{882}{13,010}\approx 6.78\%$
55 274,947 6,619 $\frac{6619}{274,947}\approx 2.41\%$
66 274,947 6,619 $\frac{6619}{274,947}\approx 2.41\%$
*Case and hospitalization projections are obtained from Table 2 of [11] by dividing, respectively, cases-prevented estimates by 0.30 and hospitalizations-prevented estimates by 0.72.
@@
$\bullet$ The FDA assumed in its most likely scenario (as well as Scenarios 5, and 6) that case and hospitalization rates would remain constant at peak 2021 Omicron levels for 5 months. Throughout the pandemic, the number of COVID infections has always been significantly larger than the corresponding number of reported cases. For the period February 2020 through September 2021, the CDC estimated the number of infections to be four times the corresponding number of reported cases [25]. Moreover, the case-to-infection multiplier likely increased (e.g. owing to increased home testing) during the Omicron-dominant period in the U.S., which started mid-December 2021. If we conservatively assume that the CDC’s 4:1 infection:case ratio holds over the evaluation period, we find the FDA’s evaluation-period case-count projection of 274,947 per 1 million unvaccinated 18–25-year-old males (Scenarios 1, 5, 6) corresponds to 4 · 274,947 = 1,099,788 infections, well over 1 million. Rounding down, we deem the FDA’s case-total assumption to be equivalent to the assumption that among one million unvaccinated males, there would be, over 5 months, 1,000,000 infections. Having well over one million infections in a group of 1 million 18–25 year-olds over 5 months is implausible. Multiple infections over a five-month period would be unlikely given the high protection from reinfection resulting from recovery from a recent infection, and any reinfections would be offset by individuals who did not become infected at all (e.g., owing to prior-infection protection, limited exposure to infectious persons, or, possibly, genetic factors [37]). [^4] The preprint [37], posted 6/10/21, was later published [38]. Even direct inoculation with SARS-CoV-2 does not necessarily produce infection: in the challenge study [39] of the 36 inoculated, 16 “remained uninfected.” Thus, for 3 of the FDA’s 5 Omicron-based scenarios, including its most likely scenario, the FDA essentially assumed that every unvaccinated male 18–25 would be infected by COVID over the evaluation period.
2.3. The FDA’s mRNA-1273 Risks Assessment

### v4

# A Reanalysis of the FDA’s Benefit–Risk Assessment of Moderna’s mRNA-1273 COVID Vaccine Based on a Model Incorporating Benefits Derived from Prior COVID Infection

@@
Abstract.
Background: The U.S. Food and Drug Administration (FDA) conducted a benefit–risk assessment for Moderna’s COVID vaccine mRNA-1273 prior to its full approval, announced 31 January 2022. The FDA’s assessment focused on males 18–64 years old because its risk analysis was limited to vaccine-attributable myocarditis/pericarditis (VAM/P), given the excess risk among males. The FDA’s analysis concluded that vaccine benefits outweighed risks, even for 18–25-year-old males (those at highest VAM/P risk). We reanalyze the FDA’s benefit–risk assessment using information available through the third week of January 2022 and focusing on 18–25-year-old males.
Methods: We develop a benefit–risk model, extending the FDA’s, that can stratify benefits and risks of vaccination by prior-infection and comorbidity status. We use the FDA’s framework but apply our model to account for benefits derived from prior COVID infection, while also accounting for finer age stratification in COVID-hospitalization rates, incidental hospitalizations (those of patients who test positive for COVID but receive treatment for something else), more realistic projections of Omicron-infection rates, and more accurate VAM/P rates.
Results: With hospitalizations as the principal endpoint of the analysis (those prevented by vaccination vs. those caused by VAM/P), our model finds vaccine risks outweighed benefits for 18–25-year-old males, except in scenarios projecting implausibly high Omicron-infection prevalence. Our assessment suggests that mRNA-1273 vaccination of 18–25-year-old males generated between 8% and 52% more hospitalizations for VAM/P compared to COVID hospitalizations prevented (over a five-month period of vaccine protection assumed by the FDA). The preceding assessment uses model inputs based on data available at the time of the FDA’s mRNA-1273 assessment. Moreover, these inputs as well as model outputs are validated by subsequently available data.
Conclusions: The outcome of a vaccine benefit–risk assessment may be dramatically impacted by accounting for the benefits derived from prior infection by the vaccine-targeted disease. To increase public confidence in vaccines and thereby reduce vaccine hesitancy, public-health agencies should employ benefit–risk models capable of supporting stratification of vaccination recommendations not only based on age and sex but also on prior-infection and comorbidity status.
@@ 1. Introduction
Because vaccines can be given to large populations of healthy persons and can be introduced as mandatory, robust benefit–risk analyses must be conducted to ensure that vaccines have highly favorable benefit–risk profiles [1].
@@
The Food and Drug Administration (FDA) issued emergency-use authorizations for mRNA COVID-19 vaccines from Pfizer (BNT162b2) and Moderna (mRNA-1273) in December 2020 [2]. Once the mRNA COVID-19 vaccines were given to large populations, it became evident that they could cause myocarditis or pericarditis (hereafter myo/pericarditis), inflammation of the heart or of the membrane surrounding the heart, respectively, with studies indicating the highest risk among males under 40 years old [3, 4, 5, 6, 7, 8, 9]. Possible mechanisms by which mRNA COVID vaccination causes myocarditis are explored in  [10, 11].
@@
The FDA published two benefit–risk analyses of primary series mRNA vaccination, the first for BNT162b2 [12] and the second for mRNA-1273 [13]. To assess vaccine benefits, the FDA estimated vaccine-preventable COVID-19 cases, hospitalizations, intensive-care unit (ICU) admissions, and deaths [12, 13]. To assess vaccine risks, the FDA estimated vaccine-attributable myo/pericarditis (VAM/P) cases, hospitalizations, ICU admissions, and deaths [12, 13]. The FDA concluded that the benefits of the mRNA vaccines outweighed VAM/P risk even for males in its highest-risk age groups: 16–17 years for BNT162b2 [12] and 18–25 years for mRNA-1273 [13]. To simplify our exposition, we focus mainly on the FDA’s analysis for Moderna’s mRNA-1273—there are significant differences in modeling assumptions in the FDA’s Pfizer and Moderna assessments. Also, VAM/P risk is substantially greater for mRNA-1273 than for BNT162b2 [3, 4, 5, 6, 7, 8, 9].
Our benefit–risk modeling differs from the FDA’s primarily by its accounting for the protection conferred by prior infection, and, at the time of the FDA’s Moderna analysis (mid-January 2022), likely over 70% of 18–25-year-olds in the U.S. had been infected by COVID-19—see Supplement S1. On the risk side, we use data available at the time of the FDA’s mRNA-1273 assessment, including that from [14, 15, 16, 12, 17], suggesting VAM/P risk for mRNA-1273 exceeded FDA estimates; see Section 1.4 below.
@@
Here, our principal aim is to demonstrate how to incorporate the protection derived from prior infection into a population-level quantitative benefit–risk model for a two-dose vaccination series. Our model has hospitalizations as its endpoint, comparing the hospitalizations-prevented benefit of vaccination to the hospitalization risk attributable to vaccination. We illustrate the utility of our model by applying it to extend and improve the FDA’s mRNA-1273 assessment relative to its hospitalizations endpoint, which can be viewed as the most important of the agency’s endpoints owing to the public-health concern that during waves of the COVID pandemic hospitals might be so overwhelmed that they would not be able to deliver high quality care to their patients.
@@
In addition to prior-infection protection, our modeling accounts for incidental hospitalizations, age-specific infection-hospitalization rates, and a more accurate assessment of VAM/P risk. We apply our model using data available to the FDA at the time of its benefit–risk analysis for mRNA-1273, assumed to be the 3rd week of January 2022 (see Section 1.2). We present evidence, using real-world data, that our model’s estimates of the hospitalizations-prevented benefit of mRNA-1273 vaccination are significantly more accurate than those the FDA obtained in [13]. We also analyze benefits and risks of vaccination for specific subpopulations such as males 18–25 having prior-infection protection and no comorbidities. Our model can be adapted to produce a benefit–risk assessment for any vaccine designed for deployment on a population-wide scale, stratified not only based on age and sex but also on prior-infection and comorbidity status.
We emphasize that the application of our model to reanalyze the FDA’s benefit–risk assessment for mRNA-1273 is for demonstration purposes and does not constitute a comprehensive re-evaluation, which would address all of the FDA’s benefit–risk endpoints (cases, hospitalizations, ICU admissions, deaths).
1.1. Framework of the FDA’s mRNA-1273 Benefit–Risk Assessment
@@
In [13], the FDA describes its benefit–risk modeling of Moderna’s mRNA-1273 for males ages 18 to 64 years. Outcomes are provided via six scenarios reflecting uncertainties in three major model inputs: COVID-19 incidence, vaccine effectiveness against cases ($VE$) as well as hospitalization ($VEH$), and vaccine-attributable myo/pericarditis (VAM/P) rate. At the time of the FDA’s mRNA-1273 analysis, the CDC was recommending a 5-month interval between primary-series completion and a booster dose. Consequently, the FDA evaluated benefits and risks over a 5-month period, which we take to be 1 January 2022–31 May 2022. The FDA assumed that all vaccine effectiveness estimates remained constant over this 5-month evaluation period. The FDA’s major modeling assumptions are summarized in Table 1 below.
Table 1. The FDA’s mRNA-1273 model-input assumptions for 18–25-year-old males [13, Table 1].
@@
Scenario COVID-19 Incidence Vaccine Effectiveness VAM/P Rate Per Million 2nd doses [13, Table 2]
@@ 6 Same as Scenario 1 Same as Scenario 1 241 cases, 207 hospitalizations
Using inputs from the first row of Table 1 and incidence data from [13, Table 3], the FDA concluded for Scenario 1, its “most likely scenario,” that vaccinating one million males ages 18–25 years with two doses would prevent, over 5 months, 82,484 COVID-19 cases, 4766 COVID-19 hospitalizations, 1144 ICU admissions, and 51 deaths due to COVID-19, while risking 128 VAM/P cases, 110 VAM/P hospitalizations, 0 ICU admissions and 0 deaths from VAM/P [13, Table 2]. Observe that the FDA does include a high VAM/P-risk scenario (6) as well as lower-benefit scenarios (2 and 3).
@@ 1.2. The Time of the FDA’s mRNA-1273 Analysis
Although the FDA does not define in [13] the time of its benefit–risk analysis for mRNA-1273, a natural choice is the 3rd week of January 2022, which concluded 10 days before the FDA’s announcement of the full approval of mRNA-1273 on 31 January 2022 [18]. This choice is consistent with the FDA’s time-of-analysis definition for its BNT162b2 assessment: the 2nd week of August 2021 [12, Section 2.1], which concluded nine days before the FDA’s announcement of the full approval of BNT162b2 on 23 August 2021 [19]. Consequently, our reanalysis of the FDA’s benefit–risk assessment [13] is based on model inputs derived from information available before 22 January 2022. Subsequent information validating our input choices will often be provided but only via footnotes. Starting in Section 3.2, we discontinue basing our modeling on information available only before 22 January 2022; in Section 3.2, our intent is to show how our model can be modified to incorporate a comorbidity (obesity) as a risk modifier.
1.3. The FDA’s mRNA-1273 Benefit Assessment
@@
where $VE$ and $VEH$ are, respectively, the effectiveness of full mRNA-1273 vaccination against cases and against hospitalization; the FDA assumed that for Omicron $VE=0.30$ and $VEH=0.72$, as noted in Table 1 above.
[…]
How our benefit assessment differs from the FDA’s:
@@
$\bullet$ We account for the protection conferred by prior infection. At the time of the FDA’s mRNA-1273 assessment, there was substantial evidence that prior infection would provide significant protection against COVID infection and hospitalization. In a study published 5 January 2021, Dan et al. found “About 95% of subjects retained immune memory at $\sim$6 months after [COVID] infection” [20]. A meta-analysis of studies published December 2020 through August 2021 demonstrated that protection from prior COVID infection “is, at least, equivalent to the protection afforded by complete vaccination of COVID-naïve populations” [21]. The CDC acknowledged in a 29 October 2021 science brief that “The immunity provided by vaccine and prior infection are both high” [22], while another CDC study found that during the Delta wave, unvaccinated persons with prior COVID infection had significantly greater protection against infection and hospitalization than did fully vaccinated, COVID-naïve persons [23]. Moreover, in December 2021, the FDA acknowledged, “The omicron variant has significantly more mutations than previous SARS-CoV-2 variants, particularly in its S-gene, the gene that encodes the virus’s spike protein”[24], suggesting that prior infection, entailing exposure to all SARS-CoV-2 proteins (e.g., the nucleoprotein), would likely yield greater protection against Omicron than would mRNA-1273 vaccination, which entails exposure only to the spike protein of the ancestral strain of SARS-CoV-2. Among the findings of a study of SARS-CoV-2 specific B and T cell memory is “CD8+ T cell responses preferentially target the nucleoprotein, highlighting the potential importance of including the nucleoprotein in future vaccines” [25]. [^1] Consistent with the preceding paragraph’s evidence, a Qatari study found prior infection provided 52.2% effective protection against symptomatic Omicron infection and 91% effective protection against “severe, critical, or fatal” Omicron, with corresponding effectiveness estimates for two doses of mRNA-1273 being 2.2% and 66.3%, respectively [26, Figure 2C, 2D] (study period 23 December 2021–21 February 2022). Findings from the meta-analyses and systematic reviews [27, 28, 29, 30] provide convincing evidence that prior infection provided greater protection against Omicron during the evaluation period than did two-dose vaccination among the COVID-naïve [Supplement S1, Section S7].
@@
Further evidence of the need to account for prior-infection protection includes (i) a CDC estimate that 54.9% of 18–49 year-olds had been infected by COVID-19 by 30 September 2021 [31], (ii) a South African study [32] released 14 December 2021 suggesting overall effectiveness of prior- infection protection against Omicron to be about 47% [Supplement S1, Section S4], and (iii) data from England [33] released 22 December 2021 suggesting that prior infection among the unvaccinated provides significant protection from Omicron hospitalization [Supplement S1, Section S5].
@@
$\bullet$ We account for incidental COVID-19 hospitalizations. At the time of its mRNA-1273 assessment, incidental-hospitalization rates were well known to be substantial. On 9 January 2022, the director of the CDC stated, “[U]p to 40 percent of the patients who are coming in with COVID-19 are coming in $\ldots$ with something else” [34]. [^2] We searched unsuccessfully for published CDC incidental-COVID-hospitalizations data. A large U.S. study suggests an incidental rate for 0–18 year-olds typically well above 50% during its study period 1 March 2020–31 January 2022 and for those 19–64 an incidental rate above 40% for 1 January 2022–31 January 2022 [35, Figure2]. On 10 January 2022, Ontario began publishing incidental rates, with rates for January 8–15 exceeding 45% [36]. There is evidence that incidental rates are especially high for the young. For example, a major study [37] found, “As of May 2021, the cumulative rate for hospitalization associated with SARS-CoV-2 infection was 35.9 per 10,000 children” and “When limiting the analysis to include only patients with COVID-19 who required therapeutic interventions, the hospitalization rate decreased 5.5-fold to 6.5 per 10,000 children,” which suggests an incidental rate of over 80% “for children” (persons 0–17-years-old). See Supplement S1 for additional incidental-hospitalization data.
$\bullet$ We account for finer age stratification in COVID-hospitalization rates. The FDA assumed that hospital admission rates were uniform for males of ages 18–45, contradicting CDC models indicating that hospitalization rates for those 30–49 are two times the rates for 18–29-year-olds [38]. In contrast, we employ an Omicron infection-hospitalization rate estimate for precisely the group of interest—males, 18–25 years old. Our estimate, based on a model developed in [39], is consistent with COVID-hospitalization data for the U.S. military [Supplement S1, Section S2].
@@
$\bullet$ We use a CDC case-to-infections multiplier to inform our projections of Omicron-infection prevalence. The FDA assumed in its most likely scenario (as well as Scenarios 5 and 6) that case and hospitalization rates would remain constant at peak 2021 Omicron levels for 5 months. Throughout the pandemic, the number of COVID infections has always been significantly larger than the corresponding number of reported cases. For the period February 2020 through September 2021, the CDC estimated the number of infections to be four times the corresponding number of reported cases [31]. Moreover, the case-to-infection multiplier likely increased (e.g., owing to increased home testing [^3] See [40]) during the Omicron-dominant period in the U.S., which started mid-December 2021. If we conservatively assume that the CDC’s 4:1 infection:case ratio holds over the evaluation period, we find the FDA’s evaluation-period case-count projection of 274,947 per 1 million unvaccinated 18–25-year-old males (Scenarios 1, 5, 6) corresponds to 4 · 274,947 = 1,099,788 infections, well over 1 million. Rounding down, we deem the FDA’s case-total assumption to be equivalent to the assumption that among one million unvaccinated males, there would be, over 5 months, 1,000,000 infections. Having well over one million infections in a group of 1 million 18–25-year-olds over 5 months is implausible. Multiple infections over a 5-month period would be unlikely given the high protection from reinfection resulting from recovery from a recent infection, and any reinfections would be offset by individuals who did not become infected at all (e.g., owing to prior-infection protection, limited exposure to infectious persons, or, possibly, genetic factors [41]). [^4] The preprint [41], posted 10 June 2021, was later published [42]. Even direct inoculation with SARS-CoV-2 does not necessarily produce infection: in the challenge study [43] of the 36 inoculated, 16 “remained uninfected.” Thus, for 3 of the FDA’s 5 Omicron-based scenarios, including its most likely scenario, the FDA essentially assumed that every unvaccinated male 18–25 would be infected by COVID over the evaluation period.
In contrast, for our most likely scenario, we project COVID incidence over the evaluation period to be twice that of the 2nd COVID wave in the U.S. (1 October 2020–28 February 2021), arriving at an estimate that 45.6% of unvaccinated, COVID-naïve 18–25 year-old males in the U.S. would be infected by Omicron over the evaluation period—see Section S1 of Supplement S1.555Based on seroprevalence and modeling information, we estimate the actual evaluation-period infected rate was less than 50% [Supplement S2, Section S2.3].
1.4. The FDA’s mRNA-1273 Risk Assessment